# Observing the Sun with the Atacama Large Aperture Submillimeter Telescope (AtLAST): Forecasting Full-disk Observations

### v1

@@ Abstract
The Atacama Large Millimeter Array (ALMA) has revolutionised the field of solar millimetre astronomy with its high angular resolution and cadence. However, interferometers often have very limited fields of view (FOV), making targeted observations of highly dynamic phenomena such of flares challenging, as they are hard to predict. A large aperture single-dish millimetre telescope with a large FOV, such as the future Atacama Large Aperture Submillimeter Telescope (AtLAST), would prove useful in observing such phenomena, as one could scan the full solar disk on shorter timescales than current millimetre facilities. We aim to explore what FOVs, detector counts, and scan strategies are suitable for AtLAST to realistically push the required full-disk scan times below 1 minute, enabling regular observations of dynamic phenomena on the Sun. Utilising the maria code, we have been able to realistically simulate solar observations with AtLAST, and thoroughly explored how instrumental properties and scanning strategies affect the full-disk observations at the different frequency bands planned for AtLAST. As input images we used realistic solar disk maps, initially constructed from publicly available full disk maps from the Solar Dynamics Observatory (SDO), and thoroughly compared to ALMA total power maps. We find the double-circle scan pattern, currently employed at ALMA for full-disk mapping to be an acceptable way of scanning the Sun with AtLAST, too. Using small to intermediately sized instruments (1000 - 50,000 detector elements), the estimated observational cadence would be less than 1 minute across AtLAST’s frequency range with a reasonable pixel spacing. Using instruments with larger FOVs (, equivalent to 1 R⊙), we find a simple circular scan to be more efficient, achieving cadences on second time scales across the frequency range, but requiring more detector elements ( 100,000). We find that a large FOV single-dish telescope such as AtLAST could provide the solar millimetre community with hitherto unachievable observations, namely full-disk observations at high cadence and adequate resolution. With cadences potentially down to seconds, such an instrument would be ideal in the study of quickly evolving phenomena, such as flares.
keywords:
[…]
@@ 1 Introduction
Solar millimetre and submillimetre radiation mostly originates in the chromosphere, the atmospheric layer in-between the relatively cool photosphere and the hot corona (Wedemeyer et al., 2016). How the corona is heated to temperatures of the order K, three orders of magnitude hotter than the photosphere, has puzzled astronomers for decades and is known as the coronal heating problem. As all energy that is emitted from the Sun’s photosphere has to pass through the chromosphere, this layer plays a role in the quest to solve the atmospheric heating problem. Imprints of heating phenomena might be observed in the chromosphere, thus constraining their contributions to heating the atmosphere. The chromosphere can be studied through spectroscopic measurements in the visible and ultra-violet (UV) bands (Rouppe van der Voort et al., 2020; Thoen Faber et al., 2025). However, these lines are formed under non-local thermodynamic equilibrium conditions, and are therefore hard to interpret (see, e.g. de la Cruz Rodríguez et al., 2016; Bastian et al., 2018). Millimetre and submillimetre radiation from the chromosphere has the advantage of being thermal continuum emission, therefore formed under local thermodynamic equilibrium, and consequently much easier to interpret (Bastian et al., 2018). Furthermore, being within the Rayleigh-Jeans limit has the advantage that the observed intensity is linearly proportional to the temperature of the emitting plasma. Additionally, the formation height for this continuum flux has the advantage of increasing with wavelength (see Wedemeyer et al., 2016, and references therein), thus making multi-frequency solar observations in the mm-regime ideal for exploring how the chromospheric plasma temperature varies as a function of height over regions of the Sun with varying activity levels such as the Quiet Sun (QS), coronal holes, and Active Regions (ARs). This is important in trying to understand how energy is transported into and through the chromosphere, which plays a role in better understanding the atmospheric heating problem.
@@
The Atacama Large Millimeter Array (ALMA, Wootten and Thompson, 2009) has for almost a decade been one of the main drivers of solar millimetre astronomy (see Bastian et al., 2022, and references therein), with its superior spectral- and angular resolution and high cadence to other facilities observing in the same frequency range (Bastian et al., 2018, 2022; Wedemeyer et al., 2016). Ongoing and future improvements to instrumentation and data processing and calibration methods will facilitate excellent solar observations in the mm-regime for decades to come.
@@ 2.1 Thermal structure and heating of the solar atmosphere
AtLAST could contribute to the study of the heating of the solar atmosphere, both by observing possible candidates for heating mechanisms, e.g. in connection with flares (see Sect. 2.2), but also through studying the thermal structure of the solar atmosphere. The millimetre regime is suited for this, as the (sub-)millimetre radiation is formed under local thermodynamic equilibrium and within the Rayleigh-Jeans limit. The observed brightness temperature of the (sub-)mm radiation is therefore linearly dependent on the local plasma temperature, thus AtLAST could function as a linear thermometer, measuring the local plasma temperature in the solar atmosphere. By observing, preferably simultaneously, in different frequency bands, AtLAST would probe different heights in the atmosphere, thus providing valuable insight into the thermal structure of the solar atmosphere.
@@ 3.2 Realistic high-resolution input maps
To explore AtLAST’s ability to observe the Sun with a range of different instrument setups and scan strategies, realistic synthetic observations have to be made, here using the maria simulator. Realistic synthetic observations of the Sun at millimetre wavelengths require input in form of realistic maps of the Sun at these wavelengths at sufficiently high resolution. The latter should be better than the AtLAST beam across the entire considered wavelength range to successfully account for the impact of the AtLAST beam on the simulated observations. Higher resolution millimetre observations of the Sun exist, for example from ALMA (Wedemeyer et al., 2016, 2020; Bastian et al., 2022), but the field of view of these observations is only of limited size, covering only a small fraction of the solar disk. To explore AtLAST’s capability for full-disk mapping we need a full-disk input image, but no full-disk millimetre observations of the Sun exist with sufficiently high resolution. We therefore have to rely on observations at other wavelengths, mainly the ultraviolet, for input maps. However, full-disk maps in the ultraviolet are not perfect representations of the millimetre Sun, as slightly different heights in the atmosphere are probed. Nevertheless, they can act as sufficiently realistic reference maps, especially if one scales them to expected temperature levels and combine maps from different UV channels together to get a map better resembling the millimetre Sun. This was done by initially querying SDO AIA (Lemen et al., 2012; Pesnell et al., 2012) maps from the 304 and 1600 Å channels at the same epoch of existing full-disk ALMA TP observations. Full-disk ALMA TP maps are available in ALMA Band 3, Band 6 and Band 7, in the frequency bands 84-116, 211-275 and 275-373 GHz respectively (Remijan et al., 2019). After convolving these UV maps to the ALMA primary beam in the respective frequency bands (, , and at their respective central frequencies), and directly comparing them, it was found that the quiet Sun (QS) regions in ALMA maps best resembled those in AIA 304 Å, while the shapes of the active regions and plages (regions of enhanced emission in active regions) looked more similar to those observed in AIA 1600 Å. We therefore construct the required millimetre full-disk maps by taking the QS regions from AIA 304 Å, masking out the active regions and replacing them with those in AIA 1600 Å multiplied with a scaling factor. Afterwards, the intensity distribution of the constructed map was scaled to follow the temperature distribution of the reference ALMA TP map, thus creating a reference input map that both resembles an actual millimetre observation of the Sun visually, but also in the distribution of the brightness temperatures. An example of such a constructed map for the ALMA Band 3 taken on 11.03.2023 (ADS/JAO.ALMA#2022.1.01544.S) are shown in Fig. 2, where the full-resolution map is shown, as well as one convolved to the beam of an ALMA TP antenna in band 3. This map is used as the base of all maria-simulations in this paper. Although only based on a single Band 3 ALMA TP map, the constructed map is used across all considered frequencies for AtLAST in Sections. 4 and 5. This is sufficiently realistic for exploring the viability of using AtLAST for high-cadence full-disk scanning. As the brightness temperatures on the Sun vary greatly across the frequency range of AtLAST, the input maps used at different frequencies were scaled separately to cover realistic temperature ranges. This was done by utilising the relation between wavelength and brightness temperature from Loukitcheva et al. (2004) and scaling it to fit the mean brightness temperatures of available full-disk ALMA TP maps.
@@
As discussed in Wedemeyer et al. (2024), a solar multi-chroic camera at AtLAST should at least have a frequency range from 90 GHz to 660 GHz, where it was argued that although going to frequencies up to 1 THz would provide higher angular resolution, adequate observing conditions might not occur as often, making these observations unreliable if one opts to observe the Sun periodically (perhaps daily) over a long time. The (theoretical) diffraction-limited angular resolution (in radians) of a single-dish telescope is a function of the observing frequency (or wavelength) and the aperture diameter, , where is the wavelength, is the frequency, and is the aperture diameter. Therefore, observing at 950 GHz instead of 660 GHz with AtLAST, which would be diffraction limited, would result in an angular resolution of 1.59”, instead of the 2.29” at 660 GHz. For completeness we consider frequencies across AtLAST’s frequency range of 30 - 950 GHz, as simulating high frequency operations at relatively poor seeing conditions might give some early insight into AtLAST’s possible performance at these frequencies.
@@ 3.4 Considered parameter space
One of the most crucial properties of an focal-plane array instrument regarding scanning extended astronomical bodies is its instantaneous FOV. This property is discussed in more detail in Sect. 4.1.1, and depends on three parameters; the number of detectors, the spacing between the detectors, and the observing frequency. Observing at higher frequencies makes the beams of each detector smaller, meaning that a more dense configuration is needed to sufficiently sample a region on the sky. Therefore it is the highest frequencies that limit our multi-chroic instrument’s observational capability, as one can reach the same instantaneous FOV much easier at lower frequencies with fewer detectors. The synthetic observations presented in Sect. 4 are therefore at AtLAST’s highest frequency of 950 GHz, as if the instrument can successfully observe the solar disk at a certain cadence at 950 GHz, it could also achieve the same cadence at any lower frequency with even fewer detectors.
For a first generation solar instrument at AtLAST we consider instruments with up to 50,000 detectors per frequency band, although more modest setups are also studied. These detectors could be spaced with detector spacings from 0.5 to 2.0 f, yielding different instantaneous FOVs. For a second generation instrument, detector counts of up to 300,000 are considered (van Kampen, 2024).
@@ 4.1 Instrumental properties
Detector/pixel count | 50,000 per band (1st gen) | 1000 | 50,000 | 100,000 |
@@ 4.1.2 Scan pattern
There are many different scan strategies implemented at millimetre facilities across the globe. At ALMA, full-disk maps of the Sun are created by scanning with a singular 12-m antenna in a double-circle pattern (White et al., 2017). As described in White et al. (2017) the double-circle pattern consists of minor circles with diameters of half the diameter of the scanned region, whose centres steadily move along a major circle of the same diameter but centred on the scanned region. An example of such a scan is depicted in Fig. LABEL:fig:scan_patterns. White et al. (2017) mention that one of the strengths of such a pattern is that it allows the instrument to have a relatively steady velocity, so the acceleration is kept to a minimum. They mention that other methods, such as a Lissajous, requires sharp turns, and therefore higher acceleration and jerk.
Lissajous scans are similar to double-circle scans in that they are also combinations of two sinusoidal functions, but of different amplitudes, periods and phases. For the Lissajous daisy pattern, the instrument boresight scans the sky in a petal-like path, thus resembling a daisy (see Fig. LABEL:fig:scan_patterns for an example of a typical Lissajous daisy with a centre offset). This scan pattern, which is utilised at the GBT (Mason, 2003; Korngut et al., 2011), has radially increasing noise levels, usually having a well-observed central region and increasingly noise-dominated outer regions due to the decreased coverage (Korngut et al., 2011). Similar to the double-circle pattern, the Lissajous daisy has the advantage of allowing the telescope to reach high scan speeds without very sharp turns that require drastic accelerations (Korngut et al., 2011).
@@ 4.2 Synthetic observations with different instrumental setups
The following sections summarise the process of simulating synthetic scans of the solar disk. Three different instrumental setups with different scan strategies are discussed in detail. These setups are selected to cover the range of different possible properties for a 1st and a 2nd generation instrument with correspondingly small to large instantaneous FOV. The setups are named Case A - C, where the small FOV setup (Case A) is discussed in Sect. 4.2.1, the intermediately sized FOV setup (Case B) in Sect. 4.2.2, and the large FOV setup (Case C) is discussed in Sect. 4.2.3.
@@ 4.2.2 Scanning strategies for intermediate-FOV instruments (Case B)
A focal plane array of only 1000 pixels per frequency band is far below the current estimates for even a first generation instrument at AtLAST, and it is therefore interesting to test a more representative instrument. According to van Kampen (2024), a first generation multi-chroic camera at AtLAST is expected to be able to host around 50,000 detector elements per frequency band. The intermediate-FOV instrument is therefore chosen to have 50,000 pixels, and to have them spaced closer together, now on a detector spacing of 1 f-. Packing the detectors closer together results in a smaller instantaneous FOV, and to counteract this a maximum observing frequency of 670 GHz is here chosen. We then only consider the three bands centred on 100, 250, and 670 GHz. The parameters of the simulated instrument are shown in Table. 1 (Case B)
The 50,000 pixels observing up to GHz at 1 spacing resulted in an instrument with a field of view of degrees, almost of a solar radius. As before, we consider a double-circle scanning strategu, however, with a relatively large FOV, few secondary circles are required for sufficiently sampling the solar disk. It was here found that having only four minor circles per rotation around the major circle was sufficient for fully sampling the whole disk. The scan pattern is here shown in Fig. 3b with its associated boresight accelerations. With such a simple scan pattern, no drastic accelerations of the telescope are required, meaning that the full scan can be completed on short time scales. It was here found that the shortest time AtLAST could complete such a scan is seconds, almost seven times as fast as the small-FOV instrument in a much tighter double-circle pattern. This shows the large improvement offered by building instrumentation that fills more of the focal plane. The resulting synthetic observations from this scan can be found in the second row of Fig. 4, where as before all major structures from the input image have been recovered.

### v2

@@ Abstract
The Atacama Large Millimeter Array (ALMA) has revolutionised the field of solar millimetre astronomy with its high angular resolution and cadence. However, interferometers often have very limited fields of view (FOV), making targeted observations of highly dynamic phenomena such of flares challenging, as they are hard to predict. A large aperture single-dish millimetre telescope with a large FOV, such as the future Atacama Large Aperture Submillimeter Telescope (AtLAST), would prove useful in observing such phenomena, as one could scan the full solar disk on shorter timescales than current millimetre facilities. We aim to explore what FOVs, detector counts, and scan strategies are suitable for AtLAST to realistically push the required full-disk scan times below 1 minute, enabling regular observations of dynamic phenomena on the Sun. Utilising the maria code, we have been able to realistically simulate solar observations with AtLAST, and thoroughly explored how instrumental properties and scanning strategies affect the full-disk observations at the different frequency bands planned for AtLAST. As input images we used realistic solar disk maps, initially constructed from publicly available full disk maps from the Solar Dynamics Observatory (SDO), and thoroughly compared to ALMA total power maps. We find the double-circle scan pattern, currently employed at ALMA for full-disk mapping to be an acceptable way of scanning the Sun with AtLAST as well. Using small to intermediately sized instruments (1000 - 50,000 detector elements), the estimated observational cadence would be less than 1 minute across AtLAST’s frequency range with a reasonable pixel spacing. Using instruments with larger FOVs (, equivalent to 1 R⊙), we find a simple circular scan to be more efficient, achieving cadences on second time scales across the frequency range, but requiring more detector elements ( 100,000). We find that a large FOV single-dish telescope such as AtLAST could provide the solar millimetre community with hitherto unachievable observations, namely full-disk observations at high cadence and adequate resolution. With cadences potentially down to seconds, such an instrument would be ideal in the study of quickly evolving phenomena, such as flares.
The Open Journal of Astrophysics
@@ 1 Introduction
Solar millimetre and submillimetre radiation mostly originates in the chromosphere, the atmospheric layer in-between the relatively cool photosphere and the hot corona (Wedemeyer et al., 2016). How the corona is heated to temperatures of the order K, three orders of magnitude hotter than the photosphere, has puzzled astronomers for decades and is known as the coronal heating problem. As all energy that is emitted from the Sun’s photosphere has to pass through the chromosphere, this layer plays a role in the quest to solve the atmospheric heating problem. Imprints of heating phenomena might be observed in the chromosphere, thus constraining their contributions to heating the atmosphere. The chromosphere can be studied through spectroscopic measurements in the visible and ultra-violet (UV) bands (Rouppe van der Voort et al., 2020; Thoen Faber et al., 2025). However, these lines are formed under non-local thermodynamic equilibrium conditions, and are therefore hard to interpret (see, e.g. de la Cruz Rodr´ıguez et al., 2016; Bastian et al., 2018). Millimetre and submillimetre radiation from the chromosphere has the advantage of being thermal continuum emission, therefore formed under local thermodynamic equilibrium, and consequently much easier to interpret (Bastian et al., 2018). Furthermore, being within the Rayleigh-Jeans limit has the advantage that the observed intensity is linearly proportional to the temperature of the emitting plasma. Additionally, the formation height for this continuum flux has the advantage of increasing with wavelength (see Wedemeyer et al., 2016, and references therein), thus making multi-frequency solar observations in the mm-regime ideal for exploring how the chromospheric plasma temperature varies as a function of height over regions of the Sun with varying activity levels such as the Quiet Sun (QS), coronal holes, and Active Regions (ARs). This is important in trying to understand how energy is transported into and through the chromosphere, which plays a role in better understanding the atmospheric heating problem.
@@
The Atacama Large Millimeter Array (ALMA, Wootten and Thompson, 2009) has for almost a decade been one of the main drivers of solar millimetre astronomy (see Bastian et al., 2022, and references therein), with its superior spectral- and angular resolution and high cadence to other facilities observing in the same frequency range (Bastian et al., 2018, 2022; Wedemeyer et al., 2016). Ongoing and future improvements to instrumentation, data processing, and calibration methods will facilitate excellent solar observations in the mm-regime for decades to come.
@@ 2.1 Thermal structure and heating of the solar atmosphere
AtLAST could contribute to the study of the heating of the solar atmosphere, both by observing possible candidates for heating mechanisms, e.g. in connection with flares (see Sect. 2.2), but also through studying the thermal structure of the solar atmosphere. The millimetre regime is suited for this, as the (sub-)millimetre radiation is formed under local thermodynamic equilibrium and within the Rayleigh-Jeans limit. The observed brightness temperature of the (sub-)mm radiation is therefore linearly dependent on the local plasma temperature, thus AtLAST could function as a linear thermometer, measuring the local plasma temperature in the solar atmosphere. By observing, preferably simultaneously, in different frequency bands, AtLAST would probe the temperature at different heights in the atmosphere, thus providing valuable insight into the thermal structure of the solar atmosphere. This would better constrain the heights at which heating occurs, and therefore yielding important information regarding what mechanisms are likely to cause said heating.
@@ 3.2 Realistic high-resolution input maps
To explore AtLAST’s ability to observe the Sun with a range of different instrument setups and scan strategies, realistic synthetic observations have to be made, here using the maria simulator. Realistic synthetic observations of the Sun at millimetre wavelengths require input in form of realistic maps of the Sun at these wavelengths at sufficiently high resolution. The latter should be better than the AtLAST beam across the entire considered wavelength range to successfully account for the impact of the AtLAST beam on the simulated observations.
Higher resolution millimetre observations of the Sun exist, for example from ALMA (Wedemeyer et al., 2016, 2020; Bastian et al., 2022), but the field of view of these observations is only of limited size, covering only a small fraction of the solar disk. To explore AtLAST’s capability for full-disk mapping we need a full-disk input image, but no full-disk millimetre observations of the Sun exist with sufficiently high resolution. We therefore have to rely on observations at other wavelengths, mainly the ultraviolet, for input maps.
@@
However, full-disk maps in the ultraviolet are not perfect representations of the millimetre Sun, as slightly different heights in the atmosphere are probed. Nevertheless, they can act as sufficiently realistic reference maps, especially if one scales them to expected temperature levels and combine maps from different UV channels together to get a map better resembling the millimetre Sun. This was done by initially querying SDO AIA (Lemen et al., 2012; Pesnell et al., 2012) maps from the 304 and 1600 Å channels at the same epoch of existing full-disk ALMA TP observations. Full-disk ALMA TP maps are available in ALMA Band 3, Band 6 and Band 7, in the frequency bands 84-116, 211-275 and 275-373 GHz respectively (Remijan et al., 2019). After convolving these UV maps to the ALMA primary beam in the respective frequency bands (, , and at their respective central frequencies), and directly comparing them, it was found that the quiet Sun (QS) regions in ALMA maps best resembled those in AIA 304 Å, while the shapes of the active regions and plages (regions of enhanced emission in active regions) looked more similar to those observed in AIA 1600 Å. We therefore construct the required millimetre full-disk maps by taking AIA 304 map, keeping the QS regions, but replacing the active regions with those in AIA 1600 Å multiplied with a scaling factor. This was done using a simple intensity mask in the 304 Å map, the threshold of which had to be adjusted slightly at different epochs to achieve the visually best reference map.
Afterwards, the intensity distribution of the constructed map was scaled to follow the temperature distribution of the reference ALMA TP map, thus creating a reference input map that both resembles an actual millimetre observation of the Sun visually, but also in the distribution of the brightness temperatures. As the brightness temperatures on the Sun vary greatly across the frequency range of AtLAST, the input maps used at different frequencies were scaled separately to cover realistic temperature ranges. This was done by utilising the relation between wavelength and brightness temperature from Loukitcheva et al. (2004) and scaling it to fit the mean brightness temperatures of available full-disk ALMA TP maps.
An example of such a constructed map for the ALMA Band 3 taken on 11.03.2023 (ADS/JAO.ALMA#2022.1.01544.S) is shown in Fig. 2, where the full-resolution map is shown, as well as one convolved to the beam of an ALMA TP antenna in band 3. This map is used as the base of all maria-simulations in this paper. Although only based on a single Band 3 ALMA TP map, the constructed map is used across all considered frequencies for AtLAST in Sections. 4 and 5. This map was chosen due to the presence of large active regions with complicated structures covering a broad range of spatial scales, making it an ideal test case for the scan’s capability to reconstruct such complicated structures. Additionally, the parts of the map outside active regions are representative of quiescent regions, i.e. “Quiet Sun”, therefore also testing the scan’s capability to capture faint structures and thus also maps at epochs closer to solar minimum.
@@
As discussed in Wedemeyer et al. (2024), a solar multi-chroic camera at AtLAST should at least have a frequency range from 90 GHz to 660 GHz, where it was argued that although going to frequencies up to 1 THz would provide higher angular resolution, adequate observing conditions might not occur as often, making these observations unreliable if one opts to observe the Sun periodically (perhaps daily) over a long time. The (theoretical) diffraction-limited angular resolution (in radians) of a single-dish telescope is a function of the observing frequency (or wavelength) and the aperture diameter, , where is the wavelength, is the frequency, and is the aperture diameter. Therefore, observing at 950 GHz instead of 660 GHz with AtLAST, which would be diffraction limited, would result in an angular resolution of 1.59, instead of the 2.29 at 660 GHz. For completeness we consider frequencies across AtLAST’s frequency range of 30 - 950 GHz, as simulating high frequency operations at relatively poor seeing conditions might give some early insight into AtLAST’s possible performance at these frequencies.
@@ 3.4 Considered parameter space
One of the most crucial properties of a focal-plane array instrument regarding scanning extended astronomical bodies is its instantaneous FOV. This property is discussed in more detail in Sect. 4.1.1, and depends on three parameters; the number of detectors, the spacing between the detectors, and the observing frequency. Observing at higher frequencies makes the beams of each detector smaller, meaning that a more dense configuration is needed to sufficiently sample a region on the sky. Therefore it is the highest frequencies that limit our multi-chroic instrument’s observational capability, as one can reach the same instantaneous FOV much easier at lower frequencies with fewer detectors.
For a first generation solar instrument at AtLAST we consider instruments with up to 50,000 detectors per frequency band, although more modest setups are also studied. For a second generation instrument, detector counts of up to 300,000 are considered (van Kampen, 2024). These detectors could be spaced with detector spacings from 0.5 to 2.0 f, yielding different instantaneous FOVs. The f-number is here the ratio between the focal length and the aperture diameter, and is the wavelength.
@@ 4.1 Instrumental properties
| Pixel count | 50,000 per band (1st gen) | 1000 | 50,000 | 100,000 |
@@ 4.1.2 Scan pattern
There are many different scan strategies implemented at millimetre facilities across the globe. At ALMA, full-disk maps of the Sun are created by scanning with a singular 12-m antenna in a double-circle pattern (White et al., 2017). As described in White et al. (2017) the double-circle pattern consists of minor circles with diameters of half the diameter of the scanned region, whose centres steadily move along a major circle of the same diameter but centred on the scanned region. An example of such a scan is depicted in Fig. 1. White et al. (2017) mention that one of the strengths of such a pattern is that it allows the instrument to have a relatively steady velocity, so the acceleration is kept to a minimum. They mention that other methods, such as a Lissajous, requires sharp turns, and therefore higher acceleration and jerk.
Lissajous scans are similar to double-circle scans in that they are also combinations of two sinusoidal functions, but of different amplitudes, periods and phases. The Lissajous daisy pattern is a special kind of Lissajous, where the instrument boresight scans the sky in a petal-like path, thus resembling a daisy. See van Marrewijk et al. (2024) for an example using a Lissajous daisy scan to mimic MUSTANG-2 (Dicker et al., 2014) observations with maria. This scan pattern, which is utilised at the GBT (Mason, 2003; Korngut et al., 2011), has radially increasing noise levels, usually having a well-observed central region and increasingly noise-dominated outer regions due to the decreased coverage (Korngut et al., 2011). Similar to the double-circle pattern, the Lissajous daisy has the advantage of allowing the telescope to reach higher scan speeds than typical Lissajous patterns, without very sharp turns that require drastic accelerations (Korngut et al., 2011).
@@ 4.2 Synthetic observations with different instrumental setups
The following sections summarise the process of simulating synthetic scans of the solar disk. Three different instrumental setups with different scan strategies are discussed in detail. These setups are selected to cover the range of different possible properties for a 1st and a 2nd generation instrument with correspondingly small to large instantaneous FOV. The setups are named Case A - C, where the small FOV setup (Case A) is discussed in Sect. 4.2.1, the intermediately sized FOV setup (Case B) in Sect. 4.2.2, and the large FOV setup (Case C) is discussed in Sect. 4.2.3. These three instrumental setups all have circular detector arrays, and are summarised in Table. 1. For setups with multiple frequency bands, separate detector arrays are setup for each of the four bands, equivalent to having separate arrays in the same focal plane or splitting the light into the considered frequencies before it reaches the individual detectors.
@@ 4.2.2 Scanning strategies for intermediate-FOV instruments (Case B)
A focal plane array of only 1000 pixels per frequency band is far below the current estimates for even a first generation instrument at AtLAST, and it is therefore interesting to test a more representative instrument. According to van Kampen (2024), a first generation multi-chroic camera at AtLAST is expected to be able to host around 50,000 detector elements per frequency band. The intermediate-FOV instrument is therefore chosen to have 50,000 pixels, and to have them spaced closer together, now on a detector spacing of 1 f. Packing the detectors closer together results in a smaller instantaneous FOV, and to counteract this a maximum observing frequency of 670 GHz is here chosen. We then only consider the three bands centred on 100, 250, and 670 GHz. The parameters of the simulated instrument are shown in Table. 1 (Case B)
The 50,000 pixels observing up to GHz at 1 spacing resulted in an instrument with a field of view of degrees, almost of a solar radius. As before, we consider a double-circle scanning strategy, however, with a relatively large FOV, few secondary circles are required for sufficiently sampling the solar disk. It was here found that having only four minor circles per rotation around the major circle was sufficient for fully sampling the whole disk. The scan pattern is here shown in Fig. 3b with its associated boresight accelerations. With such a simple scan pattern, no drastic accelerations of the telescope are required, meaning that the full scan can be completed on short time scales. It was here found that the shortest time AtLAST could complete such a scan is seconds, almost seven times as fast as the small-FOV instrument in a much tighter double-circle pattern. This shows the large improvement offered by building instrumentation that fills more of the focal plane. The resulting synthetic observations from this scan can be found in the second row of Fig. 4, where as before all major structures from the input image have been recovered.